scruffie the cat has $15 to spend each month on cat toys, which cost $3 each, and cat treats, which cost $1.50 each. draw a budget line to show the combination of each good that scuffie can afford if she spends her entire budget. now suppose that cat treats go on sale for $1 each. now does this change in price affect the budget line (show on the graph)?
Answer #1

With a budget of $15

Number of cat toys you can buy = 15/3 = 5

Number of cat treats you can buy = 15/1.5 = 10

Budget line:- BL1

Cat toys BL1 10 Cat treats

When the cat treats go for sale for $1

Number of cat treats you can buy now = 15/1 =15

so, the budget line changes to BL2

Cat toys BL2 BL1 10 15 Cat treats